Find the area of a triangle whose vertices are (0,0), (1,2), (-1,2)
Answer by Alan3354(69443)

Find the area of a triangle whose vertices are A(0,0), B(1,2), C(-1,2)
---------
(1,2) and (-1,2) have the same y value, so BC parallel to the x-axis.
The length of BC is 2.
The distance from the Origin is 2.
Area = b*h/2 = 2*2/2 = 2 sq units